# A Survey on the Application of Large Language Models in Scenario-Based Testing of Automated Driving Systems

### v1

@@ V-B LLM Usage Strategy
Prompt Engineering | Zero-shot Learning | LLMs can perform complex reasoning tasks in a zero-shot manner without any task-specific examples [144]. |
@@
Few-shot Learning | It refers to the models’ ability to perform tasks by conditioning on a few examples provided in the input prompt, without any fine-tuning [5]. | |
@@
Chain-of-Thought | It is defined as a series of intermediate natural language reasoning steps that lead to the final output [145]. | |
@@
Format Constraint | It refers to that LLMs output contents following predefined formats and relevant investigations have been introduced in. | |
@@
Parameter Tuning | Full Fine-tuning | It refers to the process of updating all parameters of a pre-trained LLM to adapt it comprehensively to a new domain [146]. |
@@
Low-Rank Adaption (LoRA) | Instead of updating all parameters, it adds small trainable layers while keeping the original model frozen, thereby reducing training cost and memory usage while maintaining strong performance on new tasks [147]. | |
@@
Knowledge- Augmented Generation | Retrieval-Augmented Generation (RAG) | It integrates a retriever to fetch relevant documents from an external corpus with a generator that conditions on both the retrieved content and the original input to produce more informed and accurate responses [148]. |
@@
Knowledge Graph Integration | Knowledge graphs are integrated to enhance LLMs’ reasoning and factual accuracy [149]. |
@@
Figure 8 presents a heatmap illustrating the distribution of LLM usage strategies (❶–❽) across scenario-based testing phases (①–⑥). The analysis indicates that early testing phases, particularly scenario generation (②), exhibit the highest concentration of LLM applications. Among the various strategies, prompt engineering is the most frequently employed, with format constraints (❹) and few-shot learning (❷) being especially prevalent. Among these, format constraint stands out as the most commonly adopted strategy overall. In contrast, parameter tuning (❺ and ❻) and knowledge-augmented generation techniques (❼ and ❽) appear less frequently and are mainly confined to scenario generation (②). Minimal engagement is observed in ADSs assessment (⑥), suggesting that the integration of LLMs into the later-stage validation process remains limited and presents opportunities for future exploration.
V-C Tasks Performed by LLM
In Section IV-B, reviewed studies are classified based on the roles LLMs play in various tasks. To offer a more detailed understanding, specific tasks undertaken by LLMs are outlined in Table I and Table II.
V-D Integrated Simulator
@@
Figure 9 illustrates the frequency of simulator usage in the reviewed studies. CARLA [101] is the most widely adopted platform, appearing in 15 studies. LGSVL [150] and Esmini [125] are each used in 7 studies, followed by SUMO [151], which is used in 5. Other simulators, including MetaDrive [152], CarMaker [153], and Isaac Gym [154], are each employed in only a single study. These results suggest a strong preference for open-source environment such as CARLA and LGSVL in the current research landscape, while the adoption of commercial remains limited.
VI Industrial perspective
@@
Table IV presents an overview of how LLMs are adopted by industry in the context of scenario-based testing for ADSs. The organizations are categorized by region, with corresponding tasks and testing phases indicated.
@@
In China, 51WORLD and Huawei adopt LLMs during the scenario generation phase for XOSC creation. In addition, Huawei applies LLMs during the scenario source phase for data labeling and retrieval. In Europe, Denotic, Luxoft, and IAV utilize LLMs in the scenario generation phase for XOSC generation, while dSPACE and Fraunhofer incorporate them into test automation workflows. Wayve employs LLMs in the scenario source phase for large-scale data generation. In the USA, Applied Intuition uses LLMs in the scenario generation phase for scenario construction.
Although numerous academic studies have explored the application of LLMs in scenario-based testing of ADSs, industrial adoption remains relatively limited. As shown by the relatively small number of documented cases in Table IV, current applications in industry are still at an early exploratory stage. Furthermore, existing use cases are primarily confined to a narrow set of tasks, such as XOSC generation and test automation, highlighting a noticeable gap between academic advancements and their translation into practical industrial deployment.
@@
China | 51WORLD | ② | OpenX Generation | [155] |
Huawei | ①② | Data Labeling Data Retrieval Scenario Generation | [156] | |
EU | Denotic | ② | OpenX Generation | [157] |
IAV | ② | OpenX Generation | [158] | |
Luxoft | ② | OpenX Generation | [159] | |
dSPACE | ⑤ | Test Automation | [160] | |
Fraunhofer | ⑤ | Test Automation | [161] | |
Wayve | ① | Data Generation | [162] | |
USA | Applied Intuition | ② | Scenario Generation | [163] |
VII Open challenges
@@
VII-A LLM Hallucination and Output Variability
A key challenge in applying LLMs to scenario-based testing is their inherent response uncertainty, primarily in the format of hallucination and output variability. Hallucination represents to the generation of outputs that are either unfaithful to the input or factually incorrect [164, 165]. This issue arises across all testing phases, including physically implausible trajectories in data enrichment [97], annotation of non-existent in data labeling [89], invalid or non-executable code in scenario generation [56], and factually incorrect judgments in safety- or legality-critical assessments due to insufficient domain knowledge [40].
Output variability refers to nondeterministic behavior of LLMs, where identical prompts may yield divergent results. This unpredictability poses particular challenges in scenario generation, where minor variations can lead to cumulative structural errors [71]. For example, GPT-4 has been observed to generate inconsistent simulation configurations from identical prompts [44].
Several strategies have been proposed to mitigate these challenges. Huang et al. [165] categorize countermeasures into three levels: data, training, and inference and identify RAG as a promising approach. Elmaaroufi et al. [54] further propose using a secondary LLM as a verification agent to assess the output consistency with the given task description. Nonetheless, robust consistency guarantees remain an open problem.
VII-B Simulation Platform Integration

### v2

@@ IV-B LLM Usage Strategy
| Prompt Engineering | ❶Zero-shot Learning | LLMs can perform complex reasoning tasks in a zero-shot manner without any task-specific examples [kojima2022large]. |
@@
| ❷Few-shot Learning | It refers to the models’ ability to perform tasks by conditioning on a few examples provided in the input prompt, without any fine-tuning [brown2020language]. | |
@@
| ❸CoT | It is defined as a series of intermediate natural language reasoning steps that lead to the final output [wei2022chain]. | |
@@
| ❹ Format Constraint | It refers to that LLMs output contents following predefined formats and relevant investigations have been introduced in. | |
@@
| Parameter Tuning | ❺ Full Fine-tuning | It refers to the process of updating all parameters of a pre-trained LLM to adapt it comprehensively to a new domain [parthasarathy2024ultimate]. |
@@
| ❻LoRA | Instead of updating all parameters, it adds small trainable layers while keeping the original model frozen, thereby reducing training cost and memory usage while maintaining strong performance on new tasks [hu2022lora]. | |
@@
| Knowledge- Augmented Generation | ❼RAG | It integrates a retriever to fetch relevant documents from an external corpus with a generator that conditions on both the retrieved content and the original input to produce more informed and accurate responses [lewis2020retrieval]. |
@@
| ❽ Knowledge Graph Integration | Knowledge graphs are integrated to enhance LLMs’ reasoning and factual accuracy [ibrahim2024survey]. |
| USA | OpenAI | GPT-4 | - | ① | ❺ | STPA | [Charalampidou.2024] | - |
| ❹ | HARA | [Nouri.2024] | - | |||||
| ❸❹ | Visual scene editing | [Wei.2024b] | - | |||||
| ❷❹ | HARA | [Nouri.2024b] | - | |||||
| ❹ | HARA | [qi2025safety] | - | |||||
| ②⑤ | ❷❸❹ | Functional scenario generation | [aasi2024generating] | CARLA | ||||
| ❷❸❼ | Scenario generation | [Lu.2024d] | SUMO | |||||
| ❷❸❹❼ | DSL scenario generation | [cai2025text2scenario] | - | |||||
| ❸ | Scenario generation | [Chang.2024] | - | |||||
| AVUnit generation | [tian2025lmmenhancedsafetycriticalscenariogeneration] | LGSVL | ||||||
| ❹ | Accident report interpretation | [Guo.2024b] | LGSVL | |||||
| ② | AVUnit generation | [tian2024llm] | LGSVL | |||||
| Driving behavior generation | [10588843] | Esmini | ||||||
| ❷❹❼ | Functional scenario interpretation SCENIC generation | [Zhang.2024b] | CARLA | |||||
| ❶❹ | XOSC generation | [Zorin.2024] | - | |||||
| Cost function generation | [Liu.2024b] | - | ||||||
| ❶ | Scenario encoder generation | [Xia.] | - | |||||
| ❷❸ | Multimodal data interpretation | [Lu.2024] | CARLA | |||||
| Scenario generation | SUMO | |||||||
| ❶❹❼ | Driving behavior generation | [Nguyen.2024] | - | |||||
| ❶❸ | Functional scenario interpretation | [Tan.2023] | - | |||||
| Scenario generation | [Li.2024c] | - | ||||||
| ❷❹ | Loss function generation | [ZiyuanZhong.2023] | - | |||||
| DSL scenario generation | [Tang.2024b] | LGSVL | ||||||
| SCENIC generation | [MiceliBarone.2023] | CARLA | ||||||
| Traffic rule interpretation JSON generation | [deng2023target] | CARLA | ||||||
| SUMO configuration generation | [Guzay.2023] | SUMO | ||||||
| ❷❸❹ | User language interpretation | [Wei.2024] | Isaac Gym Blender | |||||
| Driving trajectory generation | ||||||||
| ⑤ | ❶❹ | Code generation | [nouri2025simulation] | Esmini | ||||
| GPT-4o | - | ②⑤ | ❷❸❹ | Functional scenario generation | [aasi2024generating] | CARLA | ||
| ② | ❷ | SCENIC generation | [Miao.2024] | CARLA | ||||
| ❷❸❹ | JSON generation | [Ruan.2024] | CARLA | |||||
| ❶❹ | Accident video classification | [zhang2025languagevisionmeetroad] | - | |||||
| GPT-4o-mini | - | ② | ❸❹ | Scenario generation | [Xu.2024] | CARLA | ||
| ❶❹ | Accident video classification | [zhang2025languagevisionmeetroad] | - | |||||
| GPT-4v | - | ① | ❹❻ | Visual grounding | [chen2024automated] | - | ||
| Text-davinci-003 | - | ⑤ | ❶ | Semantic anomaly detection | [Elhafsi.2023] | CARLA | ||
| CLIP [radford2021learning] | ① | ❹ | Image retrieval | [Rigoll.2024] | - | |||
| ❶❹ | Scenario retrieval | [Sohn.2024] | - | |||||
| GPT-3 | - | ① | ❹ | STPA | [diemert2023can] | - | ||
| ④ | ❹ | Scenario realism assessment | [Wu.2024] | |||||
| GPT-3.5 | - | ② | ❶❷❸❹❼ | SCENIC generation | [elmaaroufi2024scenicnl] | CARLA | ||
| ❶❷ | Domain knowledge distillation | [Tang.2023c] | - | |||||
| ❻ | XOSC generation | [Wang.2023] | - | |||||
| ❺ | Trajectory generation | [Zhao.2024b] | - | |||||
| ① | ❷❸ | Trajectory generation | [Guo.2024] | - | ||||
| Text-Embeeding -Ada-002 | - | ① | ❸❹ | Visual scene editing | [Wei.2024b] | - | ||
| Anthropic | Claude-instant-1.2 | - | ② | ❶❷❸❹❼ | SCENIC generation | [elmaaroufi2024scenicnl] | CARLA | |
| LMSYS | Vicuna-7B-1.5 [zheng2023judging] | ① | ❶❹ | Driving video generation | [jia2023adriver] | - |
| UW Madison | LLaVA-v1.5-7b [liu2024improved] | ① | ❷ | Driving log retrieval | [knapp2024data] | - | ||
| LLaVA-v1.5-13b [liu2024improved] | ② | ❹❻ | BEV interpretation | [xu2025chatbev] | - | |||
| BakLLaVA [BakLLaVA] | ① | ❶❹ | Scenario retrieval | [Sohn.2024] | - | |||
| LLaMA [touvron2023llama] | ②⑤ | ❻ | SCENIC generation | [Yang.2023] | LGSVL | |||
| LLaMA-2 [touvron2023llama2] | ① | ❻❽ | Scenario retrieval | [Tang.2024] | LGSVL | |||
| ④ | ❹ | Scenario realism assessment | [Wu.2024] | - | ||||
| LLaMA-2-7B [touvron2023llama2] | ② | ❻ | XOSC generation | [Wang.2023] | - | |||
| USA | Meta | LLaMA 3.1 | ②⑤⑥ | ❹❼ | Road generation | [Li.2024] | SUMO | |
| ② | ❸❹ | Functional scenario generation | [mei2025llm] | MetaDrive | ||||
| LLaMA-3 [grattafiori2024llama] | ② | ❶❹ | XOSC generation | [Zorin.2024] | - | |||
| LLaMA 3-8B | ② | ❹❻ | Driving behavior generation | [Tan.2024] | - | |||
| TinyLlama [zhang2024tinyllama] | ① | ❻ | Trajectory interpretation | [Zhong.] | - | |||
| CodeLlama [roziere2023code] | ⑤ | ❶❹ | Code generation | [nouri2025simulation] | Esmini | |||
| Salesforce | BLIP-2 [li2023blip] | ① | ❶❹ | Scenario retrieval | [Sohn.2024] | - | ||
| Sentence-T5 [ni2021sentence] | ② | ❷❹❼ | SCENIC generation | [Zhang.2024b] | CARLA | |||
| CodeGemma [team2024codegemma] | ② | ❷ | SCENIC generation | [nouri2025simulation] | Esmini | |||
| Gemma-7B [team2024gemma] | ① | ❷ | Driving log retrieval | [knapp2024data] | - | |||
| Gemini 1.0 Ultra | - | ② | ❷❹❼ | DSL scenario generation | [jiang2024scenediffuser] | - | ||
| China | Alibaba | Qwen-turbo | - | ② | ❶ | Driving behavior generation | [zhou2024humansim] | CARLA SUMO |
| Baidu | Yiyan | - | ② | ❻ | XOSC generation | [Wang.2023] | - | |
| Tsinghua Uni. | GLM-4-Air [glm2024chatglm] | ② | ❸❹ | Scenario generation | [Xu.2024] | CARLA | ||
| DeepSeek | DeepSeek-Coder [guo2024deepseek] | ⑤ | ❶❹ | Code generation | [nouri2025simulation] | Esmini | ||
| DeepSeek-R1 [guo2025deepseek] | ⑤ | ❶❹ | Code generation | [nouri2025simulation] | Esmini | |||
| ② | ❷❼ | Driving behavior analysis | [mei2025seeking] | - | ||||
| DeepSeek-V3 | ② | ❷❼ | Driving behavior analysis | [mei2025seeking] | - | |||
| ByteDance | LLaVA-NeXT- Interleave [li2024llava] | ② | ❶❹ | Accident video classification Visual grounding | [zhang2025languagevisionmeetroad] | - | ||
| France | Mistral AI | Mistral-7B-Instruct | ② | ❷❹ | Scenario generation | [tian2024enhancing] | - | |
| Mistral-7B | ④ | ❹ | Scenario realism assessment | [Wu.2024] | - | |||
| Mistral-7B | ④ | ❶❹ | Code generation | [nouri2025simulation] | Esmini | |||
| NA | NA | NA | NA | ① | ❺ | STPA | [Abbaspour.2024] | - |
| NA | NA | ② | ❹ | XOSC generation | [Zhao.2024] | Esmini |
@@
Figure 10 presents a heatmap illustrating the distribution of LLM usage strategies (❶–❽) across scenario-based testing phases (①–⑥). These phases include scenario source, scenario generation, scenario database, scenario selection, test execution, and ADS assessment [UNECE2023NATM][Nalic.2020][Riedmaier.2020]. As shown in Figure 10, early testing phases, particularly scenario generation (②), exhibit the highest concentration of LLM applications. Among the various strategies, prompt engineering is the most frequently employed, with format constraints (❹) and few-shot learning (❷) being especially prevalent. Among these, format constraint stands out as the most commonly adopted strategy overall. In contrast, parameter tuning (❺ and ❻) and knowledge-augmented generation techniques (❼ and ❽) appear less frequently and are mainly confined to scenario generation (②). Minimal engagement is observed in ADSs assessment (⑥), suggesting that the integration of LLMs into the later-stage validation process remains limited and presents opportunities for future exploration. One promising direction is test report interpretation, where LLMs transform raw metrics into concise diagnostics. For example, explaining that a TTC of 1.2 s in a high-speed cut-in reflects insufficient braking margin and an unavoidable collision. This would enhance the clarity of assessments and provide more insights for engineers and regulators.
IV-C Tasks Performed by LLM
IV-D Integrated Simulator
@@
Figure 11 illustrates the frequency of simulator usage in the reviewed studies. CARLA [pmlr-v78-dosovitskiy17a] is the most widely adopted platform, appearing in 15 studies. LGSVL [9294422] and Esmini [esmini] are each used in 7 studies, followed by SUMO [krajzewicz2002sumo], which is used in 5. Other simulators, including MetaDrive [li2021metadrive], CarMaker [carmaker], and Isaac Gym [makoviychuk2021isaac], are each employed in only a single study. These results suggest a strong preference for open-source environments such as CARLA and LGSVL in the current research landscape, while the adoption of commercial remains limited.
IV-E Integrated Dataset
In LLM-driven scenario-based testing of ADS, especially for scenario generation, datasets are essential. Numerous datasets have been released, and unlike earlier summaries (e.g. [ding2023survey]), this review emphasizes major datasets from the past five years and additionally considers captions and their sizes, which are critical for developing LLM-based frameworks. Table VI summarizes the datasets used in the reviewed works across multiple aspects.
| Dataset | Year | Real | Data Information | Annotation Type | Caption Size | Ref. | |||||||
| Image | Traj. | LiDAR | Radar | HPS | 2D | 3D | Lane | Caption | |||||
| highD [krajewski2018highd] | 2018 | - | - | - | - | - | - | - | [10588843] | ||||
[…]
| AMASS [mahmood2019amass] | 2019 | - | - | - | - | - | - | - | - | [Wei.2024] | |||
[…]
| Interaction dataset [zhan2019interaction] | 2019 | - | - | - | - | - | - | - | [Li.2024c] | ||||
[…]
| CCD [bao2020uncertainty] | 2020 | - | - | - | - | - | - | - | - | [Miao.2024] | |||
[…]
| nuScenes [caesar2020nuscenes] | 2020 | - | - | - | - | - | [ZiyuanZhong.2023] | ||||||
[…]
| nuPlan [caesar2021nuplan] | 2021 | - | - | - | - | - | [xu2025chatbev] | ||||||
[…]
| CODA [li2022coda] | 2022 | - | - | - | - | - | - | - | [Lu.2024] | ||||
[…]
| Bedlam [black2023bedlam] | 2023 | - | - | - | - | - | - | - | - | [Wei.2024] | |||
[…]
| Language-to-Interaction | 2023 | - | - | - | - | - | - | - | - | 120k | [Yang.2023] | ||
[…]
| nuScenes-QA [qian2024nuscenes] | 2024 | - | - | - | - | - | - | - | 290k | [Tang.2024] | |||
[…]
| CODA-LM | 2024 | - | - | - | - | - | - | - | 41.7k | [chen2024automated] | |||
[…]
IV-E1 Fidelity and Data Information
Data are either collected from real-world or synthetically generated. Real data are realistic yet lack critical scenarios, such as accidents. Synthetic data, by contrast, can contain such scenarios, though their realism must be ensured. For each dataset, the availability of images, trajectories, LiDAR, radar and human pose shape data is examined. These data modalities enable different tasks: images support 2D object detection, LiDAR and radar facilitate 3D object detection, and human pose shape data allow modeling of pedestrian behaviors.
IV-E2 Annotation Type and Caption Size
Annotation types determine the range of applications for a dataset. Specifically, 2D annotations include bounding boxes or pixel-level masks for objects, 3D annotations provide point clouds, 3D bounding boxes, or trajectories in spatial coordinates, and lane annotations include road topology. In addition to these common annotations in automated driving dataset, captions designed for LLMs are also investigated, with caption size reported as a key attribute. Such captions are particularly important for advancing LLM-driven automated driving tasks.
This summary does not represent a systematic dataset review but outlines several noticeable patterns. Most existing datasets are primarily vehicle-focused, while datasets specifically targeting vulnerable road user behavior remain limited. In addition, only a small number of datasets incorporate NL components to facilitate LLM applications in automated driving, and these have appeared only recently.
V Industrial perspective
@@
Table VII presents an overview of how LLMs are adopted by industry in the context of scenario-based testing for ADSs. The organizations are categorized by region, with corresponding tasks and testing phases indicated.
@@
In China, 51WORLD and Huawei adopt LLMs in the scenario generation phase for XOSC generation. In addition, Huawei applies LLMs during the scenario source phase for data labeling and retrieval. In Israel, Foretellix utilizes LLMs for scenario generation, and Cognata uses LLMs for data generation. In Europe, Deontic, Luxoft, and IAV utilize LLMs in the scenario generation phase for XOSC generation, while dSPACE and Fraunhofer incorporate them into test automation workflows. Wayve employs LLMs in the scenario source phase for large-scale data generation; BeamNG applies LLMs for XODR generation. In the USA, Applied Intuition uses LLMs in the scenario generation phase for scenario construction. Waymo and Microsoft report LLM-based scenario generation capabilities, Parallel Domain applies LLMs for data labeling.
@@
| Asia | 51WORLD | ② | XOSC Generation | [aigc2025] |
| Huawei | ①② | Data Labeling Data Retrieval Scenario Generation | [huawei_octopus] | |
| Foretellix | ② | Scenario Generation | [Foretellix2025] | |
| Cognata | ① | Data Generation | [Cognata2025] | |
| EU | Deontic | ② | XOSC Generation | [deontic_ai] |
| IAV | ② | XOSC Generation | [iavProducts2025] | |
| Luxoft | ② | XOSC Generation | [luxoft_genai_autonomous] | |
| dSPACE | ⑤ | Test Automation | [dspace_aws2024] | |
| Fraunhofer | ⑤ | Test Automation | [FraunhoferIESE2025] | |
| Wayve | ① | Data Generation | [hu2023gaia] | |
| BeamNG | ② | XODR Generation | [Gambi2025RoadGPT] | |
| USA | Applied Intuition | ② | Scenario Generation | [applied2025scenario] |
| Waymo | ② | Scenario Generation | [Hawkins2024WaymoGemini] | |
| Parallel Domain | ① | Data Labeling | [Dey2023Reactor] | |
| Microsoft | ② | Scenario Generation | [Microsoft2025Scenario] |
Although numerous academic studies have explored the application of LLMs in scenario-based testing of ADSs, industrial adoption remains relatively limited. As shown by the relatively small number of documented cases in Table VII, current applications in industry are still at an early exploratory stage. Furthermore, existing use cases are primarily confined to a narrow set of tasks, such as XOSC generation and test automation, highlighting a noticeable gap between academic advancements and their translation into practical industrial deployment.
Key challenges include ensuring that generated artifacts comply with industry standards (e.g., XOSC) and align with physical constraints. Another concern is achieving reproducibility and traceability, so that each generated version can be audited. In addition, the computational cost and response latency of LLM-based generation remain limiting factors for large-scale deployment, and all outputs must conform to safety standards (e.g., ISO 21448 [ISO21448]).
Looking ahead, promising directions include using coverage metrics to guide test execution, automatically identifying and comparing failure cases across system versions, and aggregating test evidence to support safety validation. Further opportunities lie in enhancing requirements-to-test traceability and enabling continuous improvement through operational feedback. A practical pathway toward mature adoption involves three steps: first, using standardized interfaces to ensure interoperability among tools; second, establishing a basic and consistent set of evaluation scenarios and metrics as a common benchmark; third, implementing governance and auditing mechanisms to maintain safety, traceability, and regulatory compliance. Together, these measures enable a gradual transition from early experimental successes to robust mid- and late-phase capabilities.
VI Open challenges
@@
VI-A LLM Hallucination and Output Variability
A key challenge in applying LLMs to scenario-based testing is their inherent response uncertainty, primarily in the format of hallucination and output variability. Hallucination refers to the generation of outputs that are either unfaithful to the input or factually incorrect [macpherson2013hallucination, huang2025survey]. This issue arises across all testing phases, including physically implausible trajectories in data enrichment [Guo.2024], annotation of non-existent objects in data labeling [chen2024automated], invalid or non-executable code in scenario generation [Zhang.2024b], and factually incorrect judgments in safety- or legality-critical assessments due to insufficient domain knowledge [Xu2024WhatTE].
Output variability refers to nondeterministic behavior of LLMs, where identical prompts may yield divergent results. This unpredictability poses particular challenges in scenario generation, where minor variations can lead to cumulative structural errors [Tang.2023c]. For example, GPT-4 has been observed to generate inconsistent simulation configurations from identical prompts [Guzay.2023].
Several strategies have been proposed to mitigate these challenges. Huang et al. [huang2025survey] categorize countermeasures into three levels: data, training, and inference and identify RAG as a promising approach. Elmaaroufi et al. [elmaaroufi2024scenicnl] further propose using a secondary LLM as a verification agent to assess the output consistency with the given task description. Nonetheless, robust consistency guarantees remain an open problem.
Future research may focus on three directions. First, enhancing verification and constraint mechanisms, for example rule-based enforcement to assure the output content quality. Second, developing domain-specific models that integrate diverse driving scenarios and regulatory knowledge to reduce hallucination risk. Third, improving determinism through averaging multiple outputs, selecting the most consistent result, or providing confidence estimates. Progress along these directions would strengthen the reliability and trustworthiness of LLMs for scenario-based testing of ADS.
VI-B Simulation Platform Integration